Suppose today is May 1, 2016, and your firm produces breakfast cereal and needs 200,000 bushels of corn in July 2016 for an upcoming promotion. You would like to lock in your costs today because you are concerned that corn prices might rise between now and July. Each contract is for 5,000 bushels; the settle price for July 2016 is $5.18 per bushel. Suppose corn prices are $5.14 per bushel in July. What will your cumulative mark to market be? (Do not round intermediate calculations. Enter your answer as a positive value if a profit or as a negative number if a loss. Round to the nearest whole number, i.e. dollar, e.g., 32.). Cumulative MTM = $_______ Hint; To hedge the price on 200,000 bushels, you buy 40 contracts. If you buy contracts, you bet the price in July will be higher than the original. It is not, so you lost and you have a negative CumMTM.
Answer #1

Cumulative MTM = (Price per Bushel in July-Settlement Price per Bushel)*Total Volume = (5.14-5.18)*200000 = -0.04*200000 = $-8000